# Adaptive Evidence Budgeting for Scalable Long-Document Reranking with LLMs

### v4

@@ Abstract
Decoder-only LLM rerankers are powerful but often struggle with long documents: inference is costly and relevance signals can be diluted as irrelevant text accumulates in the context window. Motivated by an attention analysis showing that relevance-aligned heads degrade when non-relevant text is appended, we propose EviRerank, a scalable framework that (i) scores document blocks with a lightweight selector (BM25, bi-encoder, or cross-encoder), (ii) constructs a compact evidence context under a strict token budget, and (iii) reranks with a decoder-only LLM. Our key contribution is Adaptive Evidence Budgeting (AEB), an information-density-aware dynamic stopping strategy that avoids low-utility tail blocks, and we further study Summary Augmentation (SA) within the same budget. Across TREC DL’19, DL’23, and MLDR-zh, EviRerank consistently improves over full-document LLM reranking and strong block-selection baselines while substantially reducing the required input length. On TREC DL’19, EviRerank achieves 0.743 nDCG@10 and 0.307 MAP, improving over RankLLaMA (0.701/0.288) by +0.042 nDCG@10 (+6.0%) and +0.019 MAP (+6.6%).
Adaptive Evidence Budgeting for Scalable Long-Document Reranking with LLMs
@@ 1 Introduction
Large language models (LLMs) have become strong rerankers for web search and other retrieval settings, thanks to their improved instruction following and reasoning ability. A practical bottleneck, however, is that real-world documents are often long, while decoder-style rerankers operate under a strict input budget. When the relevant evidence is sparse and scattered, naively truncating the document or feeding a fixed prefix can miss the decisive content, leading to unstable ranking quality.
@@
A common workaround is to decompose a document into smaller blocks and rerank based on a subset of blocks. This “block-first” strategy introduces a new question: how should we select and pack evidence blocks under a fixed budget? Most existing pipelines implicitly assume a fixed packing pattern (e.g., always taking a fixed number of top blocks or always filling the budget), which can be suboptimal. Intuitively, documents differ greatly in information density: some contain one or two highly relevant blocks followed by many weak ones, while others have many moderately relevant blocks. A fixed packing rule wastes tokens on low-utility blocks for the former, and may still under-utilize useful evidence for the latter.
In this work, we propose EviRerank, a lightweight evidence-guided reranking framework that constructs a compact reranker input under a strict token budget. Given a query–document pair, we score document blocks with a local scorer (e.g., BM25, bi-encoder, or cross-encoder) and concatenate selected blocks as evidence. Crucially, we introduce Adaptive Evidence Budgeting (AEB): instead of always filling the budget, the selector can stop early when additional blocks provide low marginal utility, yielding a query-adaptive evidence length. We also study an optional Summary Augmentation (SA) module under the same total budget and isolate its effect with controlled ablations. Experiments on TREC DL’19, DL’23, and MLDR-zh show consistent gains over strong baselines and a better accuracy–efficiency trade-off. Overall, the results support the core claim: evidence selection is not only about which blocks, but also how much evidence is enough.
@@
We structure our study around five questions: RQ1 Mechanism. How does attention in LLM rerankers behave on long documents, and what does it reveal about the need for evidence construction? (Appendix B) RQ2 Cross-Dataset Effectiveness. How consistent are the gains across benchmarks and languages (DL’19, DL’23, and MLDR-zh)? RQ3 Selector Choice. How does the choice of selector (BM25 / bi-encoder / cross-encoder) affect the effectiveness–efficiency trade-off? RQ4 Summary Augmentation. Does Summary Augmentation (SA) improve effectiveness under the same budget? RQ5 Efficiency. What are the efficiency benefits (latency and training memory) of budget-aware evidence construction?
@@
  • Evidence-guided reranking under a strict budget. We formalize long-document reranking as evidence construction plus reranking, and instantiate a simple yet flexible framework (EviRerank) that works with different block scorers and rerankers.
  • Adaptive Evidence Budgeting (AEB). We propose a model-agnostic dynamic stopping rule based on normalized block relevance, adapting the doc-side token usage to document information density.
  • Summary Augmentation (SA) as a separable factor. We introduce a compact summary cue under the same total budget and evaluate it with principled controls to rule out “just more tokens” explanations.
  • Empirical results and ablations. We report consistent gains on TREC DL’19/DL’23 and MLDR-zh, and provide a $\times2$ factorial ablation disentangling AEB and SA effects.
@@ 2.1 Neural IR Selectors: Cross- and Bi-encoders
Neural IR with PLMs is commonly instantiated as (i) cross-encoders that jointly encode query–text pairs for accurate reranking Devlin et al. (2019); Nogueira and Cho (2019), (ii) bi-encoders that learn dense representations for efficient scoring or retrieval Karpukhin et al. (2020); Reimers and Gurevych (2019), and (iii) late-interaction models that balance token-level matching and efficiency Khattab and Zaharia (2020); Santhanam et al. (2022). In this work, we leverage these paradigms as lightweight local selectors (BM25/bi/cross) to score document blocks before LLM reranking.
@@ 3 EviRerank: Evidence-Guided LLM Reranking with Adaptive Budgeting
Figure 2 illustrates EviRerank, an evidence-guided reranking framework for long documents. EviRerank decomposes each document into blocks, scores blocks with a lightweight local scorer (standard IR or PLMs), and then composes a compact LLM input for a decoder reranker. Two components are optional and can be toggled independently: (i) Adaptive Evidence Budgeting (AEB) for information-density-aware early stopping, and (ii) Summary Augmentation (SA) that injects a short, query-agnostic summary cue under the same total budget.
@@ 3.1 Passage Segmentation
When reading a document for assessing its relevance with respect to a given query, humans tend to segment the document into pieces before scanning each piece for relevance (Ding et al., 2020). Inspired by this, block-based processing has become popular in IR. We adopt the CogLTX (Ding et al., 2020) block decomposition method, which assigns different costs to punctuation marks and prioritizes strong boundaries (e.g., “.” and “!”) to approximate human-like sentence segmentation. CogLTX segments each long text $D$ into a sequence of blocks $b_{1},\dots,b_{n}]$ via dynamic programming, ensuring each block length does not exceed a predefined maximum value $B$. In our implementation, B=63B{=}63, following Li et al. (2023b).
To ensure consistency with the final decoder reranker, we use the reranker’s tokenizer for block processing: the Llama2 tokenizer is used for Llama2-based rerankers and the Llama3 tokenizer is used for Llama3-based rerankers. For Chinese documents, punctuation marks are fullwidth rather than halfwidth [^1] https://en.wikipedia.org/wiki/Chinese_punctuation; we therefore incorporate Chinese punctuation marks into CogLTX to improve sentence boundary detection.
@@ 3.2 Local Block Scoring
3.3 Composing the LLM Input: Evidence Blocks with Optional Summary Cue
@@
Given a query $Q$ and a document segmented into blocks $\{b_{i}\}_{i=1}^{n}$, a local scorer (BM25/bi-encoder/cross-encoder; Sec. 3.2) produces block relevance scores $s_{i}$ and block lengths $L(b_{i})$. Let the decoder reranker token budget for document content be $B$ (e.g., $\approx$480 tokens), and let $k_{s}$ be the maximum number of blocks used by the summary cue (if enabled).
@@
Naive packing fills the token budget by appending blocks, but long documents often exhibit highly skewed information density (a few decisive blocks followed by a low-utility tail). To avoid diluting strong evidence with weak blocks, we introduce AEB, which early-stops evidence packing when the marginal relevance of newly considered blocks falls below a threshold, producing a query-adaptive input length (not necessarily exactly $B$ tokens).
@@
Let $\pi$ sort blocks by descending normalized score $\tilde{s}$: $\tilde{s}_{\pi_{1}}\geq\cdots\geq\tilde{s}_{\pi_{n}}$. We scan blocks in this order and greedily add $b_{\pi_{j}}$ into $K$ if it fits the token budget ($T+L(b_{\pi_{j}})\leq B$), and stop when the next block would exceed the budget (max tokens).
@@
Finally, we restore the original document order of blocks in $K$ and truncate the last block (if needed) to fit within $B$ tokens:
$K\;=\;\mathrm{Trunc}_{B}\!\Big(\mathrm{Sort}_{\text{index}}(K)\Big).$ (4)
@@
We build a compact summary cue from blocks. We obtain block embeddings $\{e_{i}\}_{i=1}^{n}$ (normalized), compute the centroid
@@
We compose the reranker input as
$D^{\prime}\;=\;K\;\|\;S,$ (7)
where $\|$ denotes concatenation.
@@
Rationale. AEB adapts how much evidence is consumed for each query–document pair, reducing low-utility blocks when evidence is concentrated. SA provides a compact query-agnostic cue that can complement the selected evidence blocks, especially when multiple facets are present. Both components are lightweight relative to decoder reranking: local scoring operates on short blocks, and SA can be computed from cached block embeddings with a top-$k$ selection.
@@ 4 Experimental Settings
We evaluate EviRerank on multiple long-document reranking benchmarks and compare it with strong baselines, mainly targeting RQ2RQ5 introduced in Section 1.
@@ 4.3 Experimental Design and Implementation Details
We evaluate EviRerank as an LLM-augmented reranker with fine-tuning. To ensure fair comparison with RankLLaMA (LLaMA2-7B backbone), we fine-tune EviRerank using the same backbone. EviRerank adopts local block scorers as modular components:
@@
corresponding to BM25, bi-encoder, and cross-encoder selectors. We optionally enable Adaptive Evidence Budgeting (AEB) (dynamic stopping) and Summary Augmentation (SA) under the same total budget; these are analyzed via factorial ablations in Section 6.
@@
We use PyTorch 2.4 with CUDA 11.8 on a single NVIDIA A100 (40GB). We fine-tune with LoRA Hu et al. (2021) while freezing the backbone, using pairwise hinge loss Li et al. (2023b) and AdamW with warmup–decay in FP16. Default LoRA hyperparameters are r=32r{=}32, $\alpha{=}64$, learning rate $5{\times}10^{-5}$, batch size 2, and gradient accumulation steps 8. All models train for one epoch on the same triplets. RankLLaMA uses gradient checkpointing (for longer inputs) and batch size 1; otherwise hyperparameters mirror EviRerank. All models are evaluated as rerankers over the same first-stage candidates (official runs for DL tracks; for MLDR-zh, BM25 top-$k$).
@@
Queries are truncated to 32 tokens. Unless otherwise specified, we use a fixed total reranker budget $p_{\max}{=}600$ for document-side content. In the fixed setting, we allocate 480 tokens to query-focused blocks. With SA, the remaining budget is used to append a short summary cue; without SA, the block budget can be increased up to 600 tokens for longer-input controls. With AEB, block selection can stop early under the same cap. In contrast, RankLLaMA directly consumes up to 4,096 tokens.
@@
We use language-specific bi-encoders and cross-encoders as selectors in EviRerank (Table 1); for MLDR-zh we adopt Chinese encoders. Unless otherwise specified, the same bi-encoder is used for both block selection and summary cue construction.
@@
AEB uses Min–Max normalization over block scores within each document and applies ratio-based early stopping with threshold $\rho$ (Eq. 3). We select $\rho$ on the dev set by sweeping $\{0.05,0.15,\ldots,0.65\}$ and then fix it for test; other AEB parameters follow the default setting (e.g., min/max kept blocks).
@@ 5.1 Dataset-wise Effectiveness Results
5.2 Selector Choice and Effectiveness–Efficiency Trade-off (RQ3)
@@
  • Bi-encoder provides a strong accuracy–efficiency balance when the encoder matches the language/domain (MLDR-zh). With cached block embeddings, query-time scoring is efficient.
@@ 5.4.1 Inference Speed
We measure end-to-end reranking latency under identical inference settings and report the wall-clock time to rerank 100 candidate documents per query. Figure 1 plots effectiveness (nDCG@10) against cost (latency per 100 documents; lower is better).
Observations. Full-document RankLLaMA is substantially slower (50.65s) while achieving lower effectiveness (0.701 nDCG@10) than our evidence-based reranking variants. In contrast, EviRerank (AEB+SA) achieves higher effectiveness (0.743–0.744 nDCG@10) at much lower latency (9.02–11.81s). Aggregation-only baselines (RankLLaMA-MaxP/AvgP) are faster (11.54–11.64s) but suffer large effectiveness drops (0.643–0.654), confirming that explicit evidence construction is critical for maintaining reranking quality under tight budgets.
Budget-aware evidence construction enables a favorable accuracy–latency trade-off: EviRerank improves effectiveness while reducing inference cost compared with full-context LLM reranking, and it substantially outperforms score-pooling variants at comparable latency.

### v5

# EviRerank: Adaptive Evidence Construction for Long-Document LLM Reranking

@@ Abstract
Decoder-only LLM rerankers struggle with long documents: inference is costly and relevance signals can be diluted by irrelevant context. Motivated by an attention analysis indicating a consistent degradation trend when non-relevant text is appended, we propose EviRerank, an evidence-based long-document reranking framework for decoder-only LLMs. EviRerank (i) scores document blocks with a lightweight selector (BM25, bi-encoder, or cross-encoder), (ii) constructs a compact reranking context under a hard token cap by dynamically budgeting evidence blocks with Adaptive Evidence Budgeting (AEB) and adding a global summary cue via Summary Augmentation (SA), and (iii) reranks with a decoder-only LLM. Across TREC DL’19, DL’23, and MLDR-zh, EviRerank consistently outperforms full-document LLM reranking and strong block-selection baselines while substantially reducing the required input length. On TREC DL’19, EviRerank achieves 0.743 nDCG@10 and 0.307 MAP, establishing a new best result and improving over RankLLaMA (0.701/0.288) by +0.042 nDCG@10 (+6.0%) and +0.019 MAP (+6.6%).
EviRerank: Adaptive Evidence Construction for Long-Document LLM Reranking
@@ 1 Introduction
Large language models (LLMs) Touvron et al. (2023) have become strong rerankers ((e.g., RankLLaMA Ma et al. (2024)) ) for web search and other retrieval settings, thanks to advances in modeling capacity and generalization. However, real-world documents are often long, decoder-style LLMs incur rapidly growing inference cost as input length increases. Moreover, when relevant evidence is sparse and scattered, the reranker input may fail to cover the decisive signals, leading to unstable ranking quality.
@@
A common workaround is to decompose a document into smaller blocks and rerank based on a subset Li et al. (2023b); Li and Gaussier (2022). This “block-first” strategy introduces a new question: how should we construct an evidence context under a hard token cap? Most existing block-based pipelines (e.g., KeyB Li et al. (2023b) / IDCM-style Hofstätter et al. (2021) selection) primarily address which blocks to keep, and often rely on an implicit fixed allocation rule (e.g., always taking a fixed number of top blocks or always filling the budget). However, documents differ greatly in information density: some contain one or two highly relevant blocks followed by many weak ones, while others have many moderately relevant blocks. As a result, a fixed allocation wastes tokens on low-utility tail blocks for the former, yet may still fail to allocate enough evidence for the latter.
In this work, we propose EviRerank, a lightweight evidence-guided reranking framework for decoder-only LLMs that constructs a compact reranker input under a hard maximum token cap. Given a query-document pair, EviRerank first scores document blocks with a local scorer (e.g., BM25 Robertson et al. (2009), bi-encoder, or cross-encoder), and then performs evidence construction by allocating the capped context budget to block evidence and global cues. Crucially, we introduce Adaptive Evidence Budgeting (AEB), a density-aware budgeting rule that determines how much block evidence to include: instead of always filling the cap, it halts further budget allocation when additional blocks provide low marginal utility, yielding a query-adaptive evidence length. In addition, we incorporate Summary Augmentation (SA) within the same cap to inject global document cues, and isolate its effect with controlled ablations. Experiments on TREC DL’19, DL’23, and MLDR-zh show consistent gains over strong baselines and a better accuracy-efficiency trade-off.
@@
We study five questions: RQ1 Mechanism: What does attention reveal about long-document evidence construction? (Appendix B) RQ2 Generalization: Do gains hold across DL’19, DL’23, and MLDR-zh? RQ3 Selectors: How do BM25/bi-/cross-encoder selectors affect the effectiveness-efficiency trade-off? RQ4 SA: Does Summary Augmentation improve effectiveness under the same cap? RQ5 Efficiency: What are the latency and training-memory benefits?
@@
  • Evidence construction for long-document LLM reranking. We cast long-document reranking as constructing a compact evidence context under a token cap and reranking with a decoder-only LLM (EviRerank), compatible with BM25/bi-/cross-encoder selectors.
  • Adaptive Evidence Budgeting (AEB). We introduce a model-agnostic, normalized-score ratio stopping rule that adaptively determines how much block evidence to include.
  • Summary Augmentation (SA). We add a compact summary cue within the same cap and evaluate it with controlled baselines to rule out “just more tokens” effects.
  • Results and analysis. We show consistent gains on DL’19, DL’23, and MLDR-zh, and provide a $\times2$ factorial ablation isolating the effects of AEB and SA.
@@ 2.1 Neural IR Selectors: Cross- and Bi-encoders
Neural IR with PLMs is commonly instantiated as (i) cross-encoders that jointly encode query-text pairs for accurate reranking Devlin et al. (2019); Nogueira and Cho (2019), (ii) bi-encoders that learn dense representations for efficient scoring or retrieval Karpukhin et al. (2020); Reimers and Gurevych (2019), and (iii) late-interaction models that balance token-level matching and efficiency Khattab and Zaharia (2020); Santhanam et al. (2022). In this work, we leverage these paradigms as lightweight local selectors (BM25/bi/cross) to score document blocks before LLM reranking.
@@ 3 EviRerank: Evidence-Guided LLM Reranking with Adaptive Budgeting
Fig. 2 illustrates EviRerank, an evidence-guided reranking framework for long documents. EviRerank decomposes each document into blocks, scores blocks with a lightweight local scorer, and then composes a compact LLM input for a decoder reranker. Two components are optional and can be toggled independently: (i) Adaptive Evidence Budgeting (AEB) for information-density-aware early stopping, and (ii) Summary Augmentation (SA) that injects a short, query-agnostic summary cue under the same total budget.
@@ 3.1 Passage Segmentation
We segment each document into blocks using the CogLTX decomposition method (Ding et al., 2020), which favors strong punctuation boundaries and uses dynamic programming to ensure each block length is bounded. Following Li et al. (2023b), we set the maximum block length to B=63B{=}63. For consistency with the final reranker, block segmentation uses the reranker tokenizer (Llama2/Llama3 accordingly), and we extend the punctuation set for Chinese to better detect sentence boundaries.
@@ 3.2 Local Block Scoring
3.3 Composing the LLM Input: Evidence Blocks with a Summary Cue
@@
Given a query $Q$ and a document segmented into blocks $\{b_{i}\}_{i=1}^{n}$, a local scorer (BM25/bi-encoder/cross-encoder; Sec. 3.2) produces block relevance scores $s_{i}$ and block lengths $L(b_{i})$. Let the decoder reranker token cap for document-side content be $B$. We allocate this cap between evidence blocks and the summary cue: $B=B_{K}+B_{S}$, where $B_{K}$ is the maximum budget for evidence blocks $K$, and $B_{S}$ is the maximum budget for the summary cue $S$ (with $B_{S}=0$ when SA is disabled). Let $k_{s}$ be the number of blocks used to form the summary cue.
@@
Naive packing fills the token budget by appending blocks, but long documents often exhibit highly skewed information density (a few decisive blocks followed by a low-utility tail). To avoid diluting strong evidence with weak blocks, we introduce AEB, which early-stops evidence packing when the marginal relevance of newly considered blocks falls below a threshold, producing a query-adaptive evidence length that does not necessarily fill the cap $B_{K}$.
@@
Let $\pi$ sort blocks by descending normalized score $\tilde{s}$: $\tilde{s}_{\pi_{1}}\geq\cdots\geq\tilde{s}_{\pi_{n}}$. We scan blocks in this order and greedily add $b_{\pi_{j}}$ into $K$ if it fits the evidence budget ($T+L(b_{\pi_{j}})\leq B_{K}$); otherwise, we stop (i.e., blocks are packed atomically rather than split).
@@
Finally, we restore the original document order of selected blocks in $K$. As a safety measure, we apply a hard truncation after composing the final input (Step 3) to ensure the total document-side content does not exceed $B$ tokens, accounting for tokenization and prefix tokens.
@@
We build a compact summary cue from blocks. We obtain block embeddings $\{e_{i}\}_{i=1}^{n}$ from the bi-encoder (shared with the selector by default) and cache them offline for efficiency, then compute the centroid
@@
We first compose the document-side content as
$\tilde{D}\;=\;K\;\|\;S,$ (6)
and then enforce the hard token cap by truncation:
[…]
where $K$ is constructed under the evidence budget $B_{K}$ and $S$ under the summary budget $B_{S}$, with $B=B_{K}+B_{S}$.
@@
Rationale. AEB adaptively allocates the evidence budget per query-document pair, avoiding low-utility tail blocks. SA adds a compact query-agnostic global cue that complements the selected evidence blocks. Both are lightweight: scoring operates on short blocks, and SA is a top-$k$ selection over cached block embeddings.
@@ 4 Experimental Settings
We evaluate EviRerank on multiple long-document reranking benchmarks and compare it with strong baselines, mainly targeting RQ2-RQ5 introduced in Section 1.
@@ 4.3 Experimental Design and Implementation Details
We evaluate EviRerank as an LLM-augmented reranker with fine-tuning. To ensure fair comparison with RankLLaMA (LLaMA2-7B [^1] https://huggingface.co/meta-llama/Llama-2-7b-hf backbone), we fine-tune EviRerank using the same backbone. EviRerank adopts local block scorers as modular components:
@@
corresponding to BM25, bi-encoder, and cross-encoder selectors. We by default enable Adaptive Evidence Budgeting (AEB) (dynamic stopping) and Summary Augmentation (SA).
@@
We use PyTorch 2.4 with CUDA 11.8 on a single NVIDIA A100 (40GB). We fine-tune with LoRA Hu et al. (2021) while freezing the backbone, using pairwise hinge loss Li et al. (2023b) and AdamW with warmup-decay in FP16. Default LoRA hyperparameters are r=32r{=}32, $\alpha{=}64$, learning rate $5{\times}10^{-5}$, batch size 2, and gradient accumulation steps 8. All models train for one epoch on the same triplets. RankLLaMA uses gradient checkpointing (for longer inputs) and batch size 1; otherwise hyperparameters mirror EviRerank. All models are evaluated as rerankers over the same first-stage candidates (official runs for DL tracks; for MLDR-zh, BM25 top-$k$).
@@
Queries are truncated to 32 tokens following Ma et al. (2024). Unless otherwise specified, we enforce a hard maximum cap of $p_{\max}{=}600$ document-side tokens. Under this cap, we compose the reranker context from (i) query-focused evidence blocks and (ii) a lightweight summary cue (SA), while never exceeding the cap. By default, we allocate up to 480 tokens to evidence blocks and up to 120 tokens to the summary cue (within the same $p_{\max}$ cap). With AEB, evidence blocks are allocated dynamically: selection may terminate early when additional blocks provide low marginal utility, yielding a shorter query-adaptive evidence length under the same hard cap. For controlled ablations, we remove SA and reallocate its token budget to evidence blocks, while keeping the same $p_{\max}{=}600$ cap; we also compare against a 3-block random summary cue under the same cap (Table 6). In contrast, RankLLaMA directly consumes up to 4,096 document-side tokens.
@@
We use language-specific cross-encoders and a multilingual bi-encoder and as selectors in EviRerank (Table 1); Unless otherwise specified, the same bi-encoder is used for both block selection and summary cue construction.
@@
AEB applies ratio-based early stopping with threshold $\rho$ (Eq. 3). For score normalization, we use None for BM25 and Min-Max normalization for neural selectors, computed over blocks within each document. This design reflects that BM25 scores are derived from a fixed lexical matching function and are typically more comparable across blocks for the same query, whereas neural relevance scores can vary in scale and dynamic range across queries/documents, making the ratio criterion sensitive without normalization. We select $\rho$ on the dev set by sweeping $\{0.05,0.10,0.15,\ldots,0.65\}$ and then fix it for test. We also tune the minimum number of kept blocks $m$ (Eq. 3) on the dev set by sweeping $\{2,3,4,5,6\}$, while other AEB parameters follow the default setting (e.g., maximum kept blocks).
@@ 5.1 Dataset-wise Effectiveness Results
5.2 Selector Choice and Effectiveness-Efficiency Trade-off (RQ3)
@@
  • Bi-encoder provides a strong accuracy-efficiency balance when the encoder matches the language/domain (MLDR-zh). With cached block embeddings, query-time scoring is efficient.
@@ 5.4.1 Inference Speed
We measure end-to-end reranking latency under identical inference settings and report the wall-clock time to rerank 100 candidate documents per query. Fig. 1 plots effectiveness (nDCG@10) against cost (latency per 100 documents; lower is better).
Observations. Full-document RankLLaMA is slower and less effective (50.65s, 0.701 nDCG@10) than our evidence-based variants, while EviRerank (AEB+SA) achieves higher effectiveness at much lower latency (0.743-0.744; 9.02-11.81s). Pooling variants (MaxP/AvgP) run at comparable latency (11.54-11.64s) but incur large effectiveness drops (0.643-0.654), indicating that explicit evidence construction is essential under tight budgets.